OpenStudy (anonymous):

About 2 over 5 of U.S. households own at the least one dog. What is 2 over 5 written as a decimal?

OpenStudy (anonymous):

what's 2 divided by 5?

OpenStudy (anonymous):

10

OpenStudy (anonymous):

Show me your work please

OpenStudy (anonymous):

2*5=10 theres ur work

OpenStudy (anonymous):

* This is the multiplication symbol not the division symbol. what is 2/5 ?

OpenStudy (anonymous):

well srry i meant that

OpenStudy (anonymous):

and idk wat that means

OpenStudy (anonymous):

do you know how to divide?

OpenStudy (anonymous):

Basically, \(\frac{2}{5}=\frac{4}{10}\) Right? And \(\frac{4}{10}=0.4\)

OpenStudy (anonymous):

yup y not

OpenStudy (anonymous):

oh thats right its by littel brothers question he just wanted to c how it was like lol:)

OpenStudy (anonymous):

ok i belive that keith is right